OpenStudy (anonymous):

Please come to my question and see my doubt which is in the lesson od Expressions, equations and formulae

OpenStudy (anonymous):

please wait i am drawing my ques

OpenStudy (anonymous):



OpenStudy (anonymous):

hi please help

OpenStudy (mathstudent55):

The perimeter of a quadrilateral is the sum of the lengths of the sides. P = L + L + W + W

OpenStudy (triciaal):

the perimeter means all the way around for a rectangle it is

OpenStudy (anonymous):

its not a quadrilateral, its a rectangle

OpenStudy (mathstudent55):

This can be simplified to P = 2L + 2W P = 2(L + W)

OpenStudy (mathstudent55):

Yes, it is the perimeter of a rectangle.

OpenStudy (anonymous):

P = 2(L + W) is the answer for what a) or b)

OpenStudy (triciaal):

@American.Eagle a quadrilaterial is a 4 sided figure example a rectangle

OpenStudy (anonymous):

ok

OpenStudy (triciaal):

quad means 4

OpenStudy (mathstudent55):

Your quadrilateral is also a rectangle.

OpenStudy (anonymous):

tell me clearly what is the answer for a) and what is the answer for b)

OpenStudy (mathstudent55):

a) L + L + W + W b) P = 2(L + W)

OpenStudy (triciaal):

if you want to know clearly you need to take the time to read the post

OpenStudy (triciaal):

ask questions about any post you do not understand but do not expect just to get an answer.

OpenStudy (anonymous):

@triciaal thanks but mathstudent55 already told me the answer clearly

OpenStudy (anonymous):

now i dont have to do a lot of hardwork

OpenStudy (triciaal):

yes he did in the very first response

OpenStudy (anonymous):

can i tell u something if u dont mind

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

can I

OpenStudy (anonymous):

tell yes or no

OpenStudy (triciaal):

@Nnesha